Grade 9Mechanics

A body starts falling from height H and travels distance h/2 to during last second of its motion then time of

surendra surendra
7 Years agoGrade 9

1 Answer

Arun
7 Years ago
time = t
height = h
h = 1/2 gt²
2h = gt²
time to travel the first half = (t-1)
so, 
h/2  = 1/2 g (t-1)²
h = g (t-1)²
divide both equations
2= t² / (t-1)²
taking under rot on both sides
t/t-1 = √2 = 1.414
t = 1.414 t - 1.414
t = 3.41 s